# Your Turn: At Home Turning Angle Estimation for Parkinson’s Disease Severity Assessment

### v2

@@ Abstract
People with Parkinson’s Disease (PD) often experience progressively worsening gait, including changes in how they turn around, as the disease progresses. Existing clinical rating tools are not capable of capturing hour-by-hour variations of PD symptoms, as they are confined to brief assessments within clinic settings, leaving gait performance outside these controlled environments unaccounted for. Measuring turning angles continuously and passively is a component step towards using gait characteristics as sensitive indicators of disease progression in PD. This paper presents a deep learning-based approach to automatically quantify turning angles by extracting 3D skeletons from videos and calculating the rotation of hip and knee joints. We utilise state-of-the-art human pose estimation models, Fastpose and Strided Transformer, on a total of 1386 turning video clips from 24 subjects (12 people with PD and 12 healthy control volunteers), trimmed from a PD dataset of unscripted free-living videos in a home-like setting (Turn-REMAP). We also curate a turning video dataset, Turn-H3.6M, from the public Human3.6M human pose benchmark with 3D ground truth, to further validate our method. Previous gait research has primarily taken place in clinics or laboratories evaluating scripted gait outcomes, but this work focuses on free-living home settings where complexities exist, such as baggy clothing and poor lighting. Due to difficulties in obtaining accurate ground truth data in a free-living setting, we quantise the angle into the nearest bin $45^{\circ}$ based on the manual labelling of expert clinicians. Our method achieves a turning calculation accuracy of 41.6%, a Mean Absolute Error (MAE) of 34.7°, and a weighted precision (WPrec) of 68.3% for Turn-REMAP. On Turn-H3.6M, it achieves an accuracy of 73.5%, an MAE of 18.5°, and a WPrec of 86.2%. This is the first work to explore the use of single monocular camera data to quantify turns by PD patients in a home setting. All data and models are publicly available, providing a baseline for turning parameter measurement to promote future PD gait research.
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative movement disorder, characterised by symptoms such as slowness of movement and gait dysfunction [Jankovic2008] which fluctuate across the day but progress slowly over the years [Holden2018]. Currently, treatment of PD relies on therapies which improve symptoms. There are no treatments available which modify the course of the underlying disease (so-called disease-modifying treatments, or DMTs), despite there being multiple putative DMTs showing promise in laboratory studies [Lang2013]. One reason for the slow development of DMTs is the dearth of sensitive, frequent, objective biomarkers to enhance the current gold-standard clinical rating scale [Goetz2008] to measure the progression of PD. This gold-standard clinical rating scale, the Movement Disorders Society-sponsored revision of the Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) [Goetz2008], includes subjective questionnaires concerning gait and mobility experiences, along with clinicians’ ratings of scripted activities performed by the participants. The assessments typically occur within clinical settings over short durations, offering only a ”snapshot” of symptoms which vary on an hourly basis. It also has limitations including its non-linear and discontinuous scoring system, the inter-rater variability [Post2005] and Hawthorne effect [Paradis2017] of being observed on how someone mobilises [Robles-Garcia2015, Morberg2018]. Gait and turning abnormalities are common features of PD with over half of the patients reporting difficulties with turning [Bloem2001] – when someone moves round on their axis while upright, changing the direction they face. Turning changes associated with PD include the ’en bloc’ phenomenon where upper and lower body segments turn simultaneously [Spildooren2013], a longer duration of turn, less accurate turn completion, a narrower base of support [Mellone2016] and the use of ’step turns’ rather than ’pivot turns’ [Hulbert2015]. More than 40% of daily steps are during turning [Glaister2007] and turning abnormalities can predispose to falls, thus turning parameters could potentially be used as measures predicting the time to falls in a patient with PD [Bloem2001]. Furthermore, if a fall happens during turning, it is up to 8 times more likely to result in a hip fracture [Cummings1994]. In unmedicated early-stage PD, gait parameters from turning are more sensitive to change compared to straight-ahead gait outcomes [Zampieri2010], making measuring aspects of turns potentially of specific use in clinical trials of disease-modifying interventions which typically recruit recently diagnosed patients [Stephenson2021]. People with PD turn differently when being watched by a clinician [Morgan2022PRD], so measuring turning passively in uncontrolled home settings (Figure 1) could give information about mobility not captured by face-to-face assessments in the clinic.
@@
Figure 1: Examples of free-living home scenario turning activity in REMAP [Morgan2023_REMAPOpen]. Sample frames (first row) in the living room and (second row) in a kitchen.
Being able to measure the angle of turn therefore could be very helpful in PD assessment, for use in clinical trials and clinical practice. Turning angle alone provides useful insight into the progression of the disease: people with PD take larger angles of turn when they are taking medications, compared to when they withhold their symptom-improving therapies [Conradsson2017]. Turning in gait also comprises other potential measurable elements including foot strike angle, arm swing and turn speed. Calculating the changes of angle over time could help to analyse and interpret these metrics of turning. Previous work shows that the number of turning steps and the turn duration, from unplanned and pre-planned turns, can distinguish between PD medication states (whether someone takes or withholds medication) [Morgan2022PRD]. Turning speed can be used to differentiate between healthy control and PD participants [Mancini2015]. These turning parameters correlate strongly with the MDS-UPDRS scores, showing their potential to evaluate disease severity and progression [Morgan2022PRD]. Therefore, an accurate and robust method to measure the magnitude of the turning serves as the cornerstone for building more sensitive markers of the disease.
@@
In this paper, we present a deep learning-based pipeline to estimate turning angles. We adopt state-of-the-art human pose estimation models to extract 3D human body joint coordinates from monocular RGB videos. The angle of the turn can be calculated by leveraging the orientation of the paired (left and right) hip and knee joints, which are on the frontal plane of the human body and serve as reliable indicators of the direction in which the body is facing. We apply the proposed pipeline on Turn-REMAP, a dataset of turning video clips trimmed from the unique REMAP dataset [Morgan2023_REMAPOpen, morgan2023multimodal], which includes unscripted spontaneous turning activities from passively collected home monitoring videos. To evaluate our proposed method, a retrospective analysis of the trimmed video clips by clinicians serves as the ground truth reference. As it is hard to acquire the precise degree of turning using the naked eye, we adopted a special quantisation method: different from the reference technique used by previous studies [Pham2017, shin2021quantitative], we classify turning angles into the nearest discrete $45^{\circ}$ bins.
@@ 2 Literature review
Gait Analysis for Parkinson’s Disease – Gait analysis plays an instrumental role in many clinical applications, and is studied closely in PD [di2020gait, zanardi2021gait]. With widely available open-source pose estimation models applied to movement videos collected during clinical assessments, most state-of-the-art works analyse such patient videos using deep learning models and compare their outcomes against clinicians’ annotations to establish the clinical meaning of the measured gait features.
@@ 3 Datasets
Turn-REMAP – REMAP [Morgan2023_REMAPOpen] includes PD and healthy participants engaging in actions, such as sit-to-stand transitions or walking turns within a home environment. These specific actions were recorded during free-living, undirected situations, as well as formal clinical evaluations. We present Turn-REMAP, a subset of this data comprising all its turning actions, loosely-scripted and spontaneous (see Figure 1). The video data is collected using Microsoft Kinect wall-mounted cameras installed on the ground floor (communal areas) of a test-bed house [sphere2015] which captured red-green-blue (RGB) and depth data 2-3 hours daily (during daylight hours at times when participants were at home). The acceptability of using such high-resolution video recordings for validation purposes in home settings in PD has been studied in [McNaney2022, Morgan2022JMIR]. Table 1 summarises the details of Turn-REMAP. The dataset contains 12 spousal/parent-child/friend-friend pairs (24 participants in total) living freely in this sensor-embedded smart home for five days at a time. Each pair consists of one person with PD and one person who was a healthy control volunteer (HC). This pairing was chosen to enable PD vs HC comparison, for safety reasons and also to increase the naturalistic social behaviour (particularly amongst the spousal pairs who already lived together). Of the 24 participants, five females and seven males have PD. The average age of the participants is 60.25 (PD 61.25, Control 59.25) and the average time since PD diagnosis for the person with PD is 11.3 years (range 0.5-19).
Table 1: Summary of Turn-REMAP details.
# Videos # Frames # PD # HC Avg. PD Age Avg. HC Age Avg. Age Avg. Time Since Diag.
@@
Figure 2: Quantised distribution of angles. – Angles are quantised into $45^{\circ}$ bins, e.g. a turning angle anywhere in the range $90^{\circ}\pm 22.5^{\circ}$ is labelled as $90^{\circ}$. The distribution plot on the graph is from Turn-H3.6M.
The RGB videos were watched post-hoc by medical doctors who had undertaken training in the MDS-UPDRS rating score, including gait parameter evaluation. Two clinicians watched up to 4 simultaneously captured video files at a time using ELAN software [aguera2011elan] to manually annotate the videos to the nearest millisecond to the extent possible by a human rater. A pre-prepared annotation template with controlled vocabularies in drop-down menus was used to reduce the variability in the annotations created [Morgan2021_labels]. The parameters annotated included: turning angle estimation ($90^{\circ}$-$360^{\circ}$ in $45^{\circ}$ increments, shown in Figure 2) and duration of turn (seconds: milliseconds). Our definition of a turning episode is characterised by the initiation of pelvis rotation, continuing until the completion of the movement, which differs from a turn made within a walking arc, like walking around a table. The duration of labelled data recorded by the cameras for PD and HC is 72.84 and 75.31 hours, respectively.
@@
Table 2: The number of turns, including their angles, in each type of activity in Turn-REMAP.
@@
Figure 3: A turning clip in Turn-H3.6M. – An example extracted from the Human3.6M dataset [h36m_pami].
Table 3: Details for different bins on curated turning videos dataset Turn-H3.6M.
@@ 4 Methodology
Figure 4: The workflow of our proposed pipeline. – Following monocular video as input, we apply FastPose (FP) [fang2022alphapose] to estimate joint locations and lift the 2D skeleton series into 3D using the Strided Transformer (ST) [stridedTransformer]. With 3D skeletons, we compute the turning angle as the rotation of joints after being projected onto the horizontal plane. A continuous value of the angle is determined and further quantised into the nearest $45^{\circ}$ bin, which is then compared to the clinician’s annotation.
3D human joints estimation – Our approach comprises a two-stage framework where we first detect 2D human joint locations in each frame of the video sequence and then reconstruct them in 3D space based on the spatial-temporal knowledge extracted from the temporal 2D skeletons series using a deep learning model. Another way of estimating 3D human pose from videos is to use a single deep learning model to infer the 3D coordinates from the RGB pixels directly in an end-to-end manner [pavlakos2017coarse, zhao2019semantic]. However, a more loosely coupled pipeline is chosen over end-to-end frameworks as it has been shown to achieve higher accuracy with significantly lower computational cost on almost all of the benchmarks for human pose estimation [martinez2017simple, stridedTransformer, zhang2022mixste].
@@
FastPose uses a top-down framework, which detects the human object from the frames and estimates the joint coordinates in the form of a heatmap within a bounding box. The model utilises the classical ResNet [resnet] as the image feature extraction backbone, and then uses upsampling modules [wang2018understanding] and 1D convolution to generate heatmaps to represent the probability of each pixel being a human joint. FastPose outputs a heatmap for each joint, selecting the pixel with the maximum value as the joint’s coordinate. Before feeding the video frames into FastPose, we apply standard preprocessing techniques [sun2019deep, fang2022alphapose]: rescaling, normalisation, and flip augmentation. The detected human bounding boxes are first rescaled to a uniform size of 256×196 resolution, as required by the model. Subsequently, the input is normalised by subtracting the mean pixel values for each RGB channel, which helps account for differences in brightness and contrast between frames. Additionally, we employ standard flip augmentation for both training and inference. In this process, we flip the input of FastPose to obtain a flipped output. By flipping the output back and averaging it with the original output, we derive the final prediction.
@@
Numerous architectures have been suggested to address this ill-posed problem [martinez2017simple, zhang2022mixste]. We adopt the state-of-the-art model, Strided Transformer [stridedTransformer], to map the 2D keypoints series $\mathbf{K}\in\mathbb{R^{\mathit{T*J*2}}}$ into 3D skeleton $\mathbf{S}\in\mathbb{R^{\mathit{}}}$. The Strided Transformer is a transformer-based architecture that converts 2D keypoints into 3D ground truth using the original transformer encoder [vaswani2017attention]. The output is then processed by another transformer encoder with strided convolutions, aggregating the sequence to reconstruct the 3D joints of the centre frame. Notably, the Strided Transformer introduces extra constraints to ensure temporal smoothness while simultaneously aggregating long-range information across the skeleton sequence.
@@
In ablations, we consider the shoulder, hip and knee joints, together and separately, and show that the combination of hip and knee vectors performs best.
The proposed turning angle estimation algorithm acts as a plug-in-and-play component for any 3D pose estimation model that produces 3D skeletons, providing a compatible method for future comparison. For each video clip, in terms of calculating the overall angle, it is mathematically the same as using only the first and last frame vectors, but the frame-by-frame manner could also inform us of how the velocity changes within one turning motion.

### v3

@@ Abstract
People with Parkinson’s Disease (PD) often experience progressively worsening gait, including changes in how they turn around, as the disease progresses. Existing clinical rating tools are not capable of capturing hour-by-hour variations of PD symptoms, as they are confined to brief assessments within clinic settings, leaving gait performance outside these controlled environments unaccounted for. Measuring turning angles continuously and passively is a component step towards using gait characteristics as sensitive indicators of disease progression in PD. This paper presents a deep learning-based approach to automatically quantify turning angles by extracting 3D skeletons from videos and calculating the rotation of hip and knee joints. We utilise advanced human pose estimation models, Fastpose and Strided Transformer, on a total of 1386 turning video clips from 24 subjects (12 people with PD and 12 healthy control volunteers), trimmed from a PD dataset of unscripted free-living videos in a home-like setting (Turn-REMAP). We also curate a turning video dataset, Turn-H3.6M, from the public Human3.6M human pose benchmark with 3D groundtruth, to further validate our method. Previous gait research has primarily taken place in clinics or laboratories evaluating scripted gait outcomes, but this work focuses on free-living home settings where complexities exist, such as baggy clothing and poor lighting. Due to difficulties in obtaining accurate groundtruth data in a free-living setting, we quantise the angle into the nearest bin $45^{\circ}$ based on the manual labelling of expert clinicians. Our method achieves a turning calculation accuracy of 41.6%, a Mean Absolute Error (MAE) of 34.7°, and a weighted precision (WPrec) of 68.3% for Turn-REMAP. On Turn-H3.6M, it achieves an accuracy of 73.5%, an MAE of 18.5°, and a WPrec of 86.2%. This is the first work to explore the use of single monocular camera data to quantify turns by PD patients in a home setting. All data and models are publicly available, providing a baseline for turning parameter measurement to promote future PD gait research.
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative movement disorder, characterised by symptoms such as slowness of movement and gait dysfunction [Jankovic2008] which fluctuate across the day but progress slowly over the years [Holden2018]. Currently, treatment of PD relies on therapies which improve symptoms. There are no treatments available which modify the course of the underlying disease (so-called disease-modifying treatments, or DMTs), despite there being multiple putative DMTs showing promise in laboratory studies [Lang2013]. One reason for the slow development of DMTs is the dearth of sensitive, frequent, objective biomarkers to enhance the current gold-standard clinical rating scale [Goetz2008] to measure the progression of PD. This gold-standard clinical rating scale, the Movement Disorders Society-sponsored revision of the Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) [Goetz2008], includes subjective questionnaires concerning gait and mobility experiences, along with clinicians’ ratings of scripted activities performed by the participants. The assessments typically occur within clinical settings over short durations, offering only a ”snapshot” of symptoms which vary on an hourly basis. It also has limitations including its non-linear and discontinuous scoring system, the inter-rater variability [Post2005] and Hawthorne effect [Paradis2017] of being observed on how someone mobilises [Robles-Garcia2015, Morberg2018].
Gait and turning abnormalities are common features of PD with over half of the patients reporting difficulties with turning [Bloem2001] – when someone moves round on their axis while upright, changing the direction they face. Turning changes associated with PD include the ’en bloc’ phenomenon where upper and lower body segments turn simultaneously [Spildooren2013], a longer duration of turn, less accurate turn completion, a narrower base of support [Mellone2016] and the use of ’step turns’ rather than ’pivot turns’ [Hulbert2015]. More than 40% of daily steps are during turning [Glaister2007] and turning abnormalities can predispose to falls, thus turning parameters could potentially be used as measures predicting the time to falls in a patient with PD [Bloem2001]. Furthermore, if a fall happens during turning, it is up to 8 times more likely to result in a hip fracture [Cummings1994]. In unmedicated early-stage PD, gait parameters from turning are more sensitive to change compared to straight-ahead gait outcomes [Zampieri2010], making measuring aspects of turns potentially of specific use in clinical trials of disease-modifying interventions which typically recruit recently diagnosed patients [Stephenson2021]. People with PD turn differently when being watched by a clinician [Morgan2022PRD], so measuring turning passively in uncontrolled home settings (Figure 1) could give information about mobility not captured by face-to-face assessments in the clinic.
@@
Figure 1: Examples of free-living home scenario turning activity in REMAP [Morgan2023_REMAPOpen] – Sample frames (first row) in the living room and (second row) in the kitchen.
Measuring turning angle has clinical value in Parkinson’s for several reasons. Firstly, it could detect whether someone has PD (compared to a healthy person) from in-home data, since people with PD take turns with smaller angles and show more variability in their turning angles throughout the day and week than control subjects [El-Gohary2013]. Secondly, it can be used to better understand the symptom severity in PD. Turning angles have been shown to illustrate PD medication status, since people with PD take larger angles of turn when they are taking medications compared to when they withhold them [Conradsson2017]. Turning angles could also be used to detect whether people with PD have the symptom of “freezing of gait”, since those subjects take on average smaller angles of turn and fewer large amplitude/angle turns at home than control subjects [Mancini2018]. Thirdly, detected turning angles can be used to predict fall risk (tested in healthy subjects), since the mean frequency of falls has been shown to be correlated with the angle of turn [yamaguchi2012effect]. Fourthly, knowing what angle a turn has been taken over can be vital in detecting and understanding any other measurements taken from gait/turning - for instance, comparing like for like, 180°180°180  ° turns can reveal early signs of disease progression in PD [salarian2009analyzing]. Being able to detect and isolate larger angle turns would improve precision in demonstrating specific PD gait features, for example, subjects who freeze in gait show increased step time variability and increased numbers of freezing episodes at larger turning angles [bhatt2013dynamics]. Finally, using simple logic, the knowledge of turning angle is vital to contextualise other turning parameters such as duration, step time variability and speed of turn in Parkinson’s disease – for example, knowing whether a turn was over 90°90°90  ° or 360°360°360  ° would add vital clinical information to the quantification of that turn’s duration.
@@
In this paper, we present a deep learning-based pipeline to estimate turning angles. We adopt advanced human pose estimation models to extract 3D human body joint coordinates from monocular RGB videos. The angle of the turn can be calculated by leveraging the orientation of the paired (left and right) hip and knee joints, which are on the frontal plane of the human body and serve as reliable indicators of the direction in which the body is facing. We apply the proposed pipeline on Turn-REMAP, a dataset of turning video clips trimmed from the unique REMAP dataset [Morgan2023_REMAPOpen, morgan2023multimodal], which includes unscripted spontaneous turning activities from passively collected home monitoring videos. To evaluate our proposed method, a retrospective analysis of the trimmed video clips by clinicians serves as the groundtruth reference. As it is hard to acquire the precise degree of turning using the naked eye, we adopted a special quantisation method: different from the reference technique used by previous studies [Pham2017, shin2021quantitative], we classify turning angles into the nearest discrete $45^{\circ}$ bins.
@@ 2 Literature review
Gait Analysis for Parkinson’s Disease – Gait analysis plays an instrumental role in many clinical applications, and is studied closely in PD [di2020gait, zanardi2021gait]. With widely available open-source pose estimation models applied to movement videos collected during clinical assessments, most leading approaches in the field analyse such patient videos using deep learning models and compare their outcomes against clinicians’ annotations to establish the clinical meaning of the measured gait features.
@@ 3 Datasets
Turn-REMAP – REMAP [Morgan2023_REMAPOpen] includes PD and healthy participants engaging in actions, such as sit-to-stand transitions or walking turns within a home environment. These specific actions were recorded during undirected and free-living situations, as well as formal clinical evaluations. We present Turn-REMAP, a subset of this data comprising all its turning actions, loosely-scripted and spontaneous (see Figure 1). The video data is collected using Microsoft Kinect wall-mounted cameras installed on the ground floor (communal areas) of a test-bed house [sphere2015] which captured red-green-blue (RGB) and depth data 2-3 hours daily (during daylight hours at times when participants were at home). The acceptability of using such high-resolution video recordings for validation purposes in home settings in PD has been studied in [McNaney2022, Morgan2022JMIR]. Table 1 summarises the details of Turn-REMAP. The dataset contains 12 spousal/parent-child/friend-friend pairs (24 participants in total) living freely in this sensor-embedded smart home for five days at a time. Each pair consists of one person with PD and one person who was a healthy control volunteer (C). This pairing was chosen to enable PD vs C comparison, for safety reasons and also to increase the naturalistic social behaviour (particularly amongst the spousal pairs who already lived together). Of the 24 participants, five females and seven males have PD. The average age of the participants is 60.25 (PD 61.25, Control 59.25) and the average time since PD diagnosis for the person with PD is 11.3 years (range 0.5-19). The PD participants were asked to withhold their dopaminergic medications (12 hours for short-acting levodopa-containing agents and 24 hours for long-acting therapies) and/or switch off their deep brain stimulators (DBS) so that they were in the practically-defined “off ” medications/DBS state for a limited period of hours [Morgan2023_REMAPOpen]. As the PD participants are in mild to moderate stages of the disease, freezing-of-gait is less evident in our dataset.
Table 1: Summary of Turn-REMAP details. C stands for control subjects
#Videos #Frames #PD #C Avg. PD Age Avg. C Age Avg. Age Avg. Time Since Diag.
@@
Figure 2: Quantised distribution of angles – Angles are quantised into $45^{\circ}$ bins, e.g. a turning angle anywhere in the range $90^{\circ}\pm 22.5^{\circ}$ is labelled as $90^{\circ}$. The distribution plot on the graph is from Turn-H3.6M. The groundtruth turning angle on the x-axis in this graph is calculated by the changes in body orientation, represented by two hip joints projected onto the ground plane in the 3D motion capture data.
The RGB videos were watched post-hoc by medical doctors who had undertaken training in the MDS-UPDRS rating score, including gait parameter evaluation. Two clinicians watched up to 4 simultaneously captured video files at a time using ELAN software [aguera2011elan] to manually annotate the videos to the nearest millisecond to the extent possible by a human rater. A pre-prepared annotation template with controlled vocabularies in drop-down menus was used to reduce the variability in the annotations created [Morgan2021_labels]. The parameters annotated included: turning angle estimation ($90^{\circ}$-$360^{\circ}$ in $45^{\circ}$ increments, shown in Figure 2) and duration of turn (seconds: milliseconds). Our definition of a turning episode is characterised by the initiation of pelvis rotation, continuing until the completion of the movement, which differs from a turn made within a walking arc, like walking around a table. The duration of labelled data recorded by the cameras for PD and C is 72.84 and 75.31 hours, respectively.
@@
Table 2: The number of turns, including their angles, in each type of activity in Turn-REMAP
@@
Figure 3: A turning clip in Turn-H3.6M – An example extracted from the Human3.6M dataset [h36m_pami].
Table 3: Details for different bins on curated turning videos dataset Turn-H3.6M
@@ 4 Methodology
Figure 4: The workflow of our proposed pipeline – a) Using monocular videos as input, we apply FastPose (FP) [fang2022alphapose] to estimate joint locations and lift the 2D skeleton series into 3D using the Strided Transformer (ST) [stridedTransformer]. With 3D skeletons, we compute the turning angle as the rotation of joints after being projected onto the horizontal plane. A continuous value of the angle is determined and further quantised into the nearest $45^{\circ}$ bin, which is then compared to the clinician’s annotation. b) The preprocessing steps of video frames before feeding into FP. c) The projection of the hip joints onto the ground plane as hip vector $\mathcal{H}_{t}$.
3D human joints estimation – Our approach comprises a two-stage framework where we first detect 2D human joint locations in each frame of the video sequence and then reconstruct them in 3D space based on the spatial-temporal knowledge extracted from the temporal 2D skeletons series using a deep learning model. Another way of estimating 3D human pose from videos is to use a single deep learning model to infer the 3D coordinates from the RGB pixels directly in an end-to-end manner [pavlakos2017coarse, zhao2019semantic]. However, a more loosely coupled pipeline is chosen over end-to-end frameworks as it has been shown to achieve higher accuracy with gre lower computational cost on almost all of the benchmarks for human pose estimation [martinez2017simple, stridedTransformer, zhang2022mixste].
@@
FastPose uses a top-down framework, which detects the human object from the frames and estimates the joint coordinates in the form of a heatmap within a bounding box. The model utilises the classical ResNet [resnet] as the image feature extraction backbone and then uses upsampling modules [wang2018understanding] and 1D convolution to generate heatmaps to represent the probability of each pixel being a human joint. FastPose outputs a heatmap for each joint, selecting the pixel with the maximum value as the joint’s coordinate. FastPose does not directly regress human joint positions from an image; instead, it must be used in conjunction with a human object detector. In our experiment, we employed YOLOv8 as an off-the-shelf human object detector. The bounding boxes detected by YOLOv8 were used to crop the input image, and these cropped regions were then fed into FastPose for joints estimation. Before feeding the frames into FastPose, we apply standard preprocessing techniques [sun2019deep, fang2022alphapose]: rescaling, normalisation, and flip augmentation. The detected human bounding boxes are first rescaled to a uniform size of 256×196 resolution, as required by the model. Subsequently, the input is normalised by subtracting the mean pixel values for each RGB channel, which helps account for differences in brightness and contrast between frames. Pixel values are divided by 255 to normalize them to the range between 0 and 1. Additionally, we employ standard flip augmentation for both training and inference. In this process, we flip the input of FastPose to obtain a flipped output. By flipping the output back and averaging it with the original output, we derive the final prediction.
@@
Numerous architectures have been suggested to address this ill-posed problem [martinez2017simple, zhang2022mixste]. We adopt the advanced model, Strided Transformer [stridedTransformer], to map the 2D keypoints series $\mathbf{K}\in\mathbb{R^{\mathit{T*J*2}}}$ into 3D skeleton $\mathbf{S}\in\mathbb{R^{\mathit{}}}$. The Strided Transformer is a transformer-based architecture that converts 2D keypoints into 3D groundtruth using the original transformer encoder [vaswani2017attention]. The output is then processed by another transformer encoder with strided convolutions, aggregating the sequence to reconstruct the 3D joints of the centre frame. Notably, the Strided Transformer introduces extra constraints to ensure temporal smoothness while simultaneously aggregating long-range information across the skeleton sequence.
@@
In ablations, we consider the shoulder, hip and knee joints, together and separately, and show that the combination of hip and knee vectors performs best. The proposed turning angle estimation algorithm acts as a plug-in-and-play component for any 3D pose estimation model that produces 3D skeletons, providing a compatible method for future comparison. For each video clip, in terms of calculating the overall angle, it is mathematically the same as using only the first and last frame vectors, but the frame-by-frame manner could also inform us of how the velocity changes within one turning motion. To evaluate the performance of turning angle estimation, we select two categorical metrics (accuracy, weighted precision) and one regression metric (mean absolute error) against the clinician’s annotation. These metrics enable us to assess both the performance of the discrete classification and quantify the magnitude of deviation between estimated and annotated turning angles.